OpenStudy (anonymous):

Which law says that the induced voltage is directly proportional to the magnetic flux? A. Ohm’s Law B. Faraday’s Law C. Ampere’s Law D. Gauss’ Law E. Edison’s Law

OpenStudy (anonymous):

this is b right?

OpenStudy (anonymous):

It should be faraday's law (b). But voltage is proportional to "rate of change of magnetic flux" and not simply "magnetic flux".

OpenStudy (anonymous):

thankyou I thought so

OpenStudy (anonymous):

What is the magnetic force acting on an electron if its speed is 3.0 × 106 meters/second and the direction is perpendicular to a magnetic field of 0.020 teslas? The value of q = -1.6 × 10-19 coulombs. A. F = 0 newtons B. F = -6.0 × 10-15 newtons C. F = -9.6 × 10-15 newtons D. F = -3.0 × 10-16 newtons E. F = -3.2 × 10-21 newtons

OpenStudy (souvik):

\[F=q*v \times B\]

OpenStudy (anonymous):

The magnetic force is given by F=Bqv sin\(\theta\) here θ is 90. Also put only magnitude of q.

OpenStudy (souvik):

there the angle between B(magnetic field) and the v(velocity of electron) is 90 degree.. so F=qvB

OpenStudy (anonymous):

do I need to calculate those then plug them in cause im bad at entering them into the equations

OpenStudy (anonymous):

you are given everything- B,q,v .just put them in eqn.